OpenStudy (gabylovesyou):

When the quadratic equation y = (x − 3)(2x + 7) is written in standard form, what is the value of the coefficient "b"? 2 −21 −1 1

OpenStudy (gorv):

@Gabylovesyou 1

OpenStudy (gabylovesyou):

why ?

OpenStudy (gorv):

ax^2+bx+c is standard equation (x-3)(2x-7)=2x^2+7x-6x-21 =2x^2+x-21 b=1 got it???

OpenStudy (anonymous):

dont give just ans gorv

OpenStudy (ryan123345):

Just foil it